OpenStudy (anonymous):

how do you differentiate y=x(sqrt(1+x^2))?

OpenStudy (cwrw238):

you can use the product rule and the chain rule to differentiate the term in the parentheses

OpenStudy (cwrw238):

write this part as (1 + x^2)^(0.5)

OpenStudy (anonymous):

i got \[(1+x^2)^{1/2}+(1/2)x(1+x^2)^{1/2}\]

OpenStudy (anonymous):

i'm not sure if i did it right

OpenStudy (cwrw238):

no 1/2 (1 + x^2)^ (-1/2) * 2x = x (1 + x^2)^(-1/2)

OpenStudy (cwrw238):

- thats the derivative of the term in brackets

OpenStudy (anonymous):

ohhh i see thanks

OpenStudy (cwrw238):

applying the product rule we get x^2(1 + x^2)^(-1/2) + (1 + x^2)^(1/2)

OpenStudy (cwrw238):

we can convert this to 1 fraction if you like = x^2 + 1 + x^2 1 + 2x^2 ----------- = ------------ sqrt(1 + x^2) sqrt(1 + x^2)